# Quality Diversity through Human Feedback

### v2

@@ 4 Experiments
4.1 Tasks
We describe our experimental setup across three benchmark tasks in the scope of robotics, RL, and computer vision. For all experiments, we use MAP-Elites (Mouret & Clune, 2015) as the basic QD algorithm. More specific implementation details can be found in Appendix B.
@@ 4.1 Tasks
We use the robotic arm domain derived from (Cully et al., 2015; Vassiliades & Mouret, 2018). The primary goal is to identify an inverse kinematics solution for each accessible position of the endpoint of a planar robotic arm equipped with revolute joints. The objective function is to minimize the variance of the joint angles. Correspondingly, a common way of measuring diversity is to record the endpoint’s positions in a 2D space. These metrics are computed using the forward kinematics of the planar arm, as detailed in Murray et al. (2017).
Maze navigation.
Kheperax (Grillotti & Cully, 2023) is a recently developed benchmark environment for conventional QD and QD-RL algorithms, which features a maze navigation task originally proposed in Mouret & Doncieux (2012). The goal of the task is to discover a collection of neural network policy controllers that facilitate the navigation of a robot across varying positions in a maze. A Khepera-like robot is used as the agent in this environment, which is equipped with laser sensors positioned at the robot’s facing directions for computing the distance. The maze is designed to be deceptive with immovable walls, making the navigation a challenging optimization task.
@@
The latent space illumination (LSI) task (Fontaine et al., 2021) is designed for applying QD algorithms to explore the latent space of a generative model. The initial LSI in Fontaine et al. (2021) demonstrates using QD to generate gameplay levels. Fontaine & Nikolaidis (2021) extends LSI to generate images of human faces that match specific text prompts.
@@
In this work, we introduce a new LSI pipeline as a more general black-box optimization task for text-to-image generation. We use Stable Diffusion (Rombach et al., 2022), a latent diffusion model with state-of-the-art capability to generate detailed images conditioned on text descriptions. In this task, the goal is to find high-quality and diverse images by optimizing the initial latent vectors, where the diffusion model is treated as a black-box. The objective is to match a text prompt scored by CLIP, but there is no pre-defined diversity metrics, which makes this task a more challenging and open-ended optimization problem in a practical, real-world setting.
4.2 Experimental Design and Evaluation
In our experiments, we design two specific scenarios. One where the ground truth diversity metric is available and one where a ground truth diversity metric is not available.
Tasks with ground truth diversity.
The first scenario leverages a predefined ground truth diversity metric to simulate human feedback, which is derived from the differences between solutions as measured by this metric. The primary reason for using simulated feedback is to facilitate the evaluation of QDHF and enable its comparison with other methods under consistent ground truth diversity metrics. Otherwise, measuring the diversity of solutions consistently across different methods would be challenging. This approach is applied to the robotic arm and maze navigation task, where the ground truth diversity metrics correspond to the position (x and y values) of the arm or robot in a 2D space. The “human judgment” is determined by the similarity of the positions, calculated as the L2 distance from the ground truth measurements.
To validate the effectiveness of our method, we benchmark against AURORA (Grillotti & Cully, 2022; Cully, 2019), which is an unsupervised diversity discovery method designed for QD, and standard QD using ground truth diversity metrics, which offers an oracle control representing the best possible performance of QD. We use four variants of AURORA, encompassing two dimension-reduction techniques: PCA and AE, and two training strategies: pre-trained (Pre) and incremental (Inc). For evaluation, solutions of each method are additionally stored in a separate archive corresponding to the ground truth diversity metrics, which is used exclusively for evaluation. We report QD score (Pugh et al., 2015) (sum of objective values, Eq. 1) and coverage (ratio of filled cells to total cells). The QD score is normalized by the archive size to a 0-100 scale for clarity.
The evaluation is conducted from two settings: 1) solutions contained within the final archive, and 2) solutions discovered by the algorithm throughout its entire search process. Specifically, the first setting evaluates the alignment of learned diversity metrics with the ground truth metrics. The second setting provides insights into the overall efficacy of the search process regardless of how the diversity is measured and maintained. This distinction is essential given QD’s fundamental premise of leveraging diversity to enhance optimization.
Tasks without ground truth diversity.
In the second scenario, there is no ground truth diversity metric and real human feedback data is used, which applies to the LSI task. To facilitate the scalability of our method, instead of using human labelers in the loop, we alternatively utilize the real human feedback data to train a preference model, and use the preference model to source feedback for training and evaluation. Similar approaches have been widely used in recent RLHF applications such as LLM training (Stiennon et al., 2020; Ouyang et al., 2022), where a preference model is trained on human feedback data, and used to predict rewards for fine-tuning the LLM.
We use the DreamSim as the preference model on image similarity, which is trained on human judgment data from the NIGHTS dataset, both from Fu et al. (2023). For comparisons, we implement a best-of-n approach as the baseline, which generates in total 2,000 images (the same number as those explored by QDHF) randomly from uniform $\mathcal{U}(0,1)$, and select a solution set of 400 images (the same number of the solution archive in QDHF) with the top CLIP scores. We also propose a heuristic as a stronger baseline that also considers diversity, which works as follows: Starting from the second step, we sample 100 latents randomly, and choose the one with a maximal l2 distance to the previous sample. This approach increases the diversity of the latents.
We conducted LSI experiments using 6 common text prompts to show generality. We provide both quantitative and qualitative results for a comprehensive evaluation. Quantitatively, we report the average CLIP score for assessing the objective, i.e., how well the generated images match the prompt. We also use DreamSim to calculate the pairwise distance between images in the solution to measure diversity. The mean pairwise distance indicates the average separation of images, and the standard deviation indicates the variability in how the images are distributed. We also conduct a user experience study ($n=46$) to assess their opinion on 1) which result is more preferred, and 2) which result is more diverse, as qualitative indicators. We display samples of solutions for qualitative assessment.
4.3 Results
We present our main results below, with more detailed results available in Appendix C.
@@ 4.3 Results
Table 1: Results for robotic arm. We report the QD score and coverage for both ”All Solutions” (solutions found throughout training) and ”Archive Solutions” (solutions in the final archive). QDHF significantly outperforms all other methods in both QD score and coverage, and closely approaches the search capability of QD using ground truth diversity metrics when considering all solutions.
@@
Robotic arm.
For the robotic arm task, detailed results are presented in Table 1. The statistics are accumulated over 20 repeated trials. QDHF significantly surpasses AURORA in terms of both QD score and coverage. The results highlight QDHF’s capability to enhance QD algorithms over unsupervised diversity discovery methods. QDHF also outperforms QDHF-baseline, a variant that separates the processes of diversity learning and QD optimization. Such results validate our hypothesis that diversity representations need to be refined during the optimization process, as finding more high-quality solutions changes the distribution of the solutions in the search space. Moreover, it is worth highlighting that, in terms of all solutions, QDHF closely matches the performance of QD with a ground truth metric. This indicates that QDHF is a potent alternative for QD in situations where manually designed metrics are not available for optimization.
Table 2: Results for maze navigation. QDHF also shows superior performance compared to other methods, and a resemblance of the search capability of QD with ground-truth diversity.
@@
QDHF-Baseline $35.6\pm 0.6$ $37.9\pm 1.1$ $\mathbf{23.7\pm 0.9}$ $24.4\pm 1.1$
@@
Results for the maze navigation task are shown in Table 2. The statistics are accumulated over 10 repeated trials because maze navigation is more computationally expensive. Similar to robotic arm, we observe that QDHF is the best for most metrics, surpassing AURORA and closely matching the performance of standard QD. The results further support the validity of QDHF as an alternative to standard QD with improved flexibility and comparable performance.
Table 3: Results for LSI. We report the CLIP score and DreamSim pairwise distance (mean and std.) as quantitative metrics. QDHF demonstrates comparable quality to Best-of-N measured by CLIP score, and significantly better diversity measured by pariwise distance. In a blind user experience study with 46 participants, QDHF outperformed Best-of-N by a considerable margin in terms of user preference and perceived diversity, with participants unaware of the origin of each output. ($*{}^{*}$ualitative indicators are ratios, and participants can select “hard to decide” in their responses.)
Quantitative Metrics Qualitative $*{}^{*}$
Method CLIP Score Mean PD Std. PD User Preference Perceived Diversity
Best-of-N $68.85$ $0.420$ $0.106$ $26.4\%$ $8.7\%$
Best-of-N+ $68.90$ $0.419$ $0.105$
QDHF $\mathbf{69.08}$ $\mathbf{0.527}$ $\mathbf{0.151}$ $\mathbf{55.1\%}$ $\mathbf{75.7\%}$
Figure 1: Qualitative result for LSI. The optimization target prompt is “an image of a bear in a national park”. The left 4x4 grid displays images with the highest CLIP scores from randomly generated images, i.e., the Best-of-N baseline. The right grid displays a uniformly-sampled subset of QDHF solutions. Qualitatively, images generated by QDHF have more variations, and show visible trends of diversity such as object sizes (large to small along x-axis) and landscape types (rocky to verdant along y-axis, terrestrial to aquatic along x-axis). The selected example closely aligns with the average user preference ratio observed in our user experience study, where QDHF results are about twice as preferred as the baseline.
@@
Results for the LSI task are shown in Table 3. The results are summarized over 6 different text prompts, and we include more details in Appendix C.2. Baseline is the best-of-n approach, and Baseline+ is the best-of-n with enhanced diversity in sampling (Sec.  4.2). Quantitatively, QDHF has a similar CLIP score to both baseline methods, but much higher mean and standard deviation of pairwise distance, which shows that QDHF is able to generate more diverse solutions while maintain the high-quality. We also observe that the Baseline+ heuristic does not produce better diversity, which indicates that there exists strong inductive bias in the diffusion model, and generating diverse images is a challenging latent optimization problem.
@@
We also compare QDHF and Baseline qualitatively through a user experience study with 46 participants, where QDHF also outperforms baseline with a considerable margin on both the user preference ratio and user diversity perception ratio. Notably, we find that most users think QDHF generates more diverse images, and a majority of them also prefers to have such diversity in the solution. An example of the LSI results is depicted in Fig 1. We can observe that images generated by QDHF have more variations, and show visible trends of diversity.
4.4 Analysis and Discussion
Scalability of QDHF.
In the context of the scalability, determining the necessary sample size of human feedback for desired performance is critical. To investigate this, we first perform an ablation study with varying sample sizes of simulated human judgments on the robotic arm task, and the results are shown in Fig. 2. The left plot demonstrates the relationship between QD score and sample size. It is evident that there is a strong correlation between the QD score and sample size. Diving deeper into this relationship, we evaluate how the latent projection captures diversity and mirrors human judgment. We calculate the accuracy of using the latents to predict the human judgment on a validation set. The right plot reveals a strong correlation between validation accuracy and the QD score, which suggests that the performance of QDHF heavily relies on the efficacy of learning the latent projection. In other words, by evaluating the judgment prediction accuracy during training, we can estimate whether the current labels are enough for QDHF to perform well.
@@ 4.4 Analysis and Discussion
Figure 2: Analysis of varying human feedback sample sizes on robotic arm. “Judgement prediction val acc” is the accuracy of the latent projection in predicting human preferences based on a validation set. There is a direct correlation between QD score and sample size, with QDHF’s performance closely tied to the accuracy of latent projection in reflecting human judgment.
@@
Alignment between learned and ground truth diversity metrics.
We evaluate the alignment of diversity metrics derived by QDHF with the underlying ground truth diversity metrics. In Fig. 3, the archives of QDHF and AURORA-Inc (PCA) are visualized for the maze navigation task. Both AURORA and QDHF appear to effectively learn a diversity space reflective of the ground truth. However, QDHF exhibits enhanced capability in discerning the relative distances between solutions, especially in under-explored areas. This suggests that while AURORA and QDHF both exhibit robust exploration capabilities within their learned diversity spaces, QDHF consistently identifies more diverse solutions. This efficacy stems from QDHF’s ability to better align its learned diversity space with the ground truth diversity, especially concerning the scales on each axis. A similar analysis for the robotic arm task is included in Appendix C.1, where QDHF also shows its advantage in modeling the scales of diversity representations for enhanced discovery of novel solutions.
Figure 3: Visualization of the solution archives on different diversity spaces for maze navigation. “GT” stands for ground truth. Each point on the heatmap is a solution with its objective value visualized in color. QDHF fills up the archives with more solutions than AURORA. While both AURORA and QDHF learned a rotated version of the maze as diversity (first column), QDHF is able to more accurately learn the scale of the maze especially in the under-explored area.

### v3

@@ 5 Experiments
5.1 Tasks and Benchmarks
We describe our experimental setup across three benchmark tasks in the domains of robotics, RL, and generative modeling. For all experiments, we use MAP-Elites (Mouret & Clune, 2015) as the QD algorithm for fair comparison. More implementation details can be found in Appendix B.
@@ 5.1 Tasks and Benchmarks
We use the robotic arm domain from Cully et al. (2015); Vassiliades & Mouret (2018). The goal is to identify an inverse kinematics solution for each accessible position of a planar robotic arm. The objective function is to minimize the variance of the joint angles. The standard way of measuring diversity is the endpoint’s positions of the arm in the 2D space, which is computed using the forward kinematics of the arm, as detailed in Murray et al. (2017).
Maze navigation (RL).
We adopt the Kheperax (Grillotti & Cully, 2023) RL environment, which features a maze navigation task originally proposed in Mouret & Doncieux (2012). The goal is to discover a collection of neural network policy that controls the agent to navigate across varying positions in a maze. A Khepera-like robot is used as the agent, which is equipped with laser sensors positioned at the robot’s facing directions for computing the distance. The maze is designed to be deceptive with immovable walls, making navigation a challenging optimization task.
@@
The latent space illumination (LSI) task (Fontaine et al., 2021) is designed for exploring the latent space of a generative model. LSI initially aims to generate different gameplay levels and has later been extended to generate images of human faces that correspond to specific text prompts as diversity (Fontaine & Nikolaidis, 2021). We introduce a new LSI benchmark on text-to-image generation with Stable Diffusion (Rombach et al., 2022), a latent diffusion model with state-of-the-art capability. In this task, the goal is to find high-quality (to match a text prompt, scored by CLIP (Radford et al., 2021)) and diverse images by optimizing the latent vectors, where the diffusion model is treated as a black box. There are no pre-defined diversity metrics, making this task a more challenging and open-ended optimization problem in a real-world setting.
This benchmark also features two types of text prompts as subtasks: singular and compositional prompts. Singular prompts elicit responses about a single concept, usually taking the form of “an image of a category” where category is drawn from labels in popular image datasets such as COCO (Lin et al., 2014) and ImageNet (Russakovsky et al., 2015). We also include a few popular text prompts used in prior work (Rombach et al., 2022) that have a similar form. For singular prompts, diversity gives more variation in images that have the same semantics, which can better meet the diverse preferences of users.
Compositional prompts are complex conjunctions of multiple concepts. We use the MCC-250 benchmark (Bellagente et al., 2023), which consists of prompts describing two objects with respective attributes, e.g., “a red apple and a yellow banana”. Recent work has shown that diffusion models often fail to correctly compose the specific objects and attributes (Chefer et al., 2023; Bellagente et al., 2023), but enhanced diversity may explore different ways of composing, and thus improve the chance of generating the correct image given a limited budget of responses.
@@
Table 1: Results for robotic arm. We report the QD score (normalized to a scale of 0-100) and coverage for “All Solutions” (solutions found throughout training) and “Archive Solutions” (solutions in the final archive). QDHF significantly outperforms QDHF-Base and AURORA, and closely approaches the search capability of QD using ground truth diversity metrics when considering all solutions.
@@
5.2 Experimental Design and Evaluation
In our experiments, we design two distinct scenarios: structured tasks, where ground truth diversity metrics are available, and open-ended tasks, where such metrics are not.
Structured tasks.
The first scenario leverages a predefined ground truth diversity metric to simulate human feedback. The primary reason for using simulated feedback is to facilitate evaluation and comparisons, which requires measuring diversity consistently across different methods. The robotic arm and maze navigation tasks fall under this category, where the ground truth diversity metrics correspond to the position (x and y values) of the arm and of the robot in the 2D space, respectively. The “human judgment” is determined by the similarity of the positions, calculated as the L2 distance from the ground truth measurements.
To validate the effectiveness of QDHF, we benchmark against AURORA (Grillotti & Cully, 2022b; Cully, 2019) and standard QD. The standard QD uses ground truth diversity metrics, which offers an oracle control representing the best possible performance of QD. For comprehensiveness, we implement four variants of AURORA, encompassing two dimension-reduction techniques: PCA and AE, and two training strategies: pre-trained (Pre) and incremental (Inc).
For evaluation, solutions of each method are additionally stored in a separate archive corresponding to the ground truth diversity metrics. We report QD score (Pugh et al., 2015) (sum of objective values, Eq. 1) and coverage (ratio of filled cells to total cells), which are standard metrics for assessing both quality and diversity of solutions. The evaluation is conducted in two settings: 1) solutions in the final archive, and 2) solutions discovered by the algorithm throughout its entire search process. The first setting evaluates the alignment of learned diversity metrics with the ground truth metrics, and the second setting provides insights into the overall efficacy of the search process regardless of how the diversity is measured and maintained.
Open-ended tasks.
In the second scenario, there is no ground truth diversity metric and real human feedback data is used in training and evaluation, which applies to the LSI task. To facilitate the efficiency and scalability of our method, instead of having human labelers in the loop, we train a preference model with real human feedback data and use the preference model to source feedback for training and evaluation. Similar approaches have been widely used in recent RLHF applications (Stiennon et al., 2020; Ouyang et al., 2022).
We use human judgment data from the NIGHTS dataset, and the DreamSim model to estimate image similarity, both from Fu et al. (2023). For comparisons, we implement a best-of-n approach on Stable Diffusion as the baseline, which generates the same number of images as QDHF with random latents, and select a solution set with top CLIP scores. We also propose a stronger baseline (Best-of-n+) featuring a heuristic that increases the diversity of the latents by sampling the next latent to be distant from previous ones.
The LSI experiments are conducted using 100 singular and 100 compositional prompts. Quantitatively, we report the average CLIP score for assessing how well the generated images match the prompt. We also use DreamSim to calculate the pairwise distance between images in the solution to measure diversity. The mean pairwise distance indicates the average separation of images, and the standard deviation indicates the variability in how the images are distributed. We also conduct human evaluation ($n=50$) using Amazon Mechanical Turk to qualitatively evaluate the results, and show examples of solutions for qualitative assessment.
Table 2: Results for maze navigation. QDHF shows superior performance compared to QDHF-Base and AURORA, as well as a resemblance of the search capability of QD with ground-truth diversity.
@@ 5.2 Experimental Design and Evaluation
QDHF-Base $35.6\pm 0.6$ $37.9\pm 1.1$ $\mathbf{23.7\pm 0.9}$ $24.4\pm 1.1$
@@
Table 3: Quantitative Results for LSI (singular prompts). We report the CLIP score and DreamSim pairwise distance (mean and std.) as quantitative metrics. QDHF demonstrates competitive quality to Stable Diffusion (Best-of-n) measured by CLIP score, and significantly better diversity measured by pariwise distance.
Method CLIP Score Mean PD Std. PD
SD (Best-of-n) $67.64$ $0.405$ $0.98$
SD (Best-of-n+) $67.87$ $0.412$ $0.102$
QDHF $\mathbf{68.47}$ $\mathbf{0.517}$ $\mathbf{0.153}$
Table 4: Human evaluation results for LSI. In a blind user experience study, QDHF outperforms Stable Diffusion (best-of-n) by considerable margins in terms of user preference and perceived diversity for singular prompts and shows a clear advantage in correctness by human judgment for compositional prompts.
Singular Prompts Compositional Prompts
Method User Preference Perceived Diversity Judgment on Correctness
Cannot Determine $9.7\%$ $8.6\%$ $6.4\%$
SD (Best-of-n) $31.9\%$ $24.8\%$ $40.6\%$
QDHF $\mathbf{58.4\%}$ $\mathbf{66.7\%}$ $\mathbf{52.8\%}$
Figure 2: Qualitative result for LSI (singular prompts). The target prompt is “an image of a bear in a national park”. The left 4x4 grid displays Stable Diffusion (Best-of-n) results, i.e., randomly generated images with the highest CLIP scores. The right grid displays a uniformly sampled subset of QDHF solutions. Qualitatively, images generated by QDHF have more variations and show visible trends of diversity such as object sizes (large to small along the x-axis) and landscape types (rocky to verdant along the y-axis, terrestrial to aquatic along the x-axis). The selected example represents the average user preference ratio observed in our user experience study, where QDHF results are about twice as preferred and three times considered as more diverse.
Figure 3: Qualitative result for LSI (compositional prompts). The target prompt is “a brown car and a red clock”. The left 3x3 grid displays SD (Best-of-n) results, where we can observe the issue of attribute leakage, i.e., the color of the clock (red) leaks to the car. However, QDHF (on the right) can generate conceptually correct responses by having more variation in the responses.
5.3 Results
We present our major results and findings below, with more detailed results available in Appendix C.
Robotic arm.
For the robotic arm task, detailed results are presented in Table 1. The statistics are accumulated over 20 repeated trials. QDHF significantly surpasses AURORA in terms of both QD score and coverage. The results highlight QDHF’s capability to enhance QD algorithms over unsupervised diversity discovery methods. QDHF also outperforms QDHF-Base, a variant that separates the processes of diversity learning and QD optimization. Such results validate our hypothesis that diversity representations need to be refined during optimization, as finding more high-quality solutions changes the distribution of the solutions in the search space. It is worth highlighting that QDHF closely matches the performance of QD with the ground truth metric when evaluating all solutions, which indicates that QDHF is a competitive alternative for QD in situations where manually designed metrics are not available.
@@ 5.3 Results
Results are shown in Table 2. The statistics are accumulated over 10 repeated trials since maze navigation is more computationally expensive. Similar to the robotic arm task, we observe that QDHF surpasses QDHF-Base and AURORA, and closely matches the performance of standard QD. The results further support the validity of QDHF as an alternative to standard QD with improved flexibility and competitive performance.
@@
Results are shown in Table 3. The results are summarized over 100 text prompts, and we include more details in Appendix C.2. SD (Best-of-n) is the best-of-n responses produced by Stable Diffusion, and SD (Best-of-n+) is the best-of-n with a heuristic to enhance diversity in sampling (Sec.  5.2). Quantitatively, QDHF has a similar CLIP score to both baseline methods, but much higher mean and standard deviation of pairwise distance, which indicates that QDHF is able to generate more diverse solutions while maintaining the high-quality. We also observe that SD (Best-of-n+) does not produce better diversity, which indicates that there exists a strong inductive bias in the diffusion model and that generating diverse images is a challenging latent optimization problem.
@@
We also compare QDHF and SD (Best-of-n) with human evaluation in a blind user study. For singular prompts, QDHF outperforms SD (Best-of-n) with a considerable margin on both the user preference ratio and user diversity perception ratio. Notably, we find that most users think QDHF generates more diverse images, and a majority of them also prefer to have such diversity in the solution. An example of the LSI results is depicted in Fig. 2, where QDHF results have more variations and show visible trends of diversity.
For compositional prompts, users are asked to select which 3x3 set of generated images most accurately responds to the prompt. We find that QDHF is more accurate in generating images with correct concepts as evaluated by humans. Fig. 3 shows an example, where QDHF mitigates the common attribute leakage issue in compositional image generation by improving the diversity of responses. It is worth highlighting that QDHF serves as a flexible black-box optimization method, compatible with any existing generative models, such as Stable Diffusion, without requiring access to or fine-tuning the model parameters.
@@
Figure 4: Analysis of varying human feedback sample sizes on robotic arm. “Judgment prediction val acc” is the accuracy of the latent projection in predicting human preferences based on a validation set. There is a direct correlation between QD score and sample size, with QDHF’s performance closely tied to the accuracy of latent projection in reflecting human judgment.
5.4 Ablation Studies and Analysis
We highlight the analysis of QDHF’s scalability and quality of learned diversity metrics. Additional analysis on robustness is provided in Appendix C.1.
Scalability.
To ensure scalability, determining the necessary amount of human feedback to achieve sufficient performance is critical. To investigate this, we perform a study of varying sample sizes of human judgments, shown in Fig. 4. The left plot demonstrates the relationship between QD score and sample size, showing a strong correlation. Diving deeper into this relationship, we evaluate how accurately the latent projection captures the diversity and mirrors human judgment using the accuracy of predicting human judgment on a validation set. The right plot reveals a strong correlation between accuracy and QD score, which suggests that QDHF relies on learning useful diversity metrics. In other words, by evaluating the accuracy of judgment prediction during training, we can estimate whether current feedback data is sufficient for QDHF to perform well.
@@ 5.4 Ablation Studies and Analysis
Alignment between learned and ground truth diversity.
We evaluate the alignment between diversity metrics derived by QDHF and the underlying ground truth diversity metrics. In Fig. 5, the archives of QDHF and AURORA-Inc (PCA) are visualized for maze navigation. Both AURORA and QDHF appear to effectively learn a diversity space reflective of the ground truth. However, QDHF exhibits an enhanced capability to discern the relative distances between solutions, especially in under-explored areas. This indicates QDHF’s strength in identifying novel solutions, and this efficacy stems from QDHF’s ability to better align its learned diversity space to the ground truth diversity, especially concerning the scales. A similar analysis for the robotic arm task is included in Appendix C.1, where QDHF also shows its advantage in modeling the scales of diversity representations for enhanced discovery of novel solutions.